### v1

# The Complexity of Maximizing the MST-ratio

Abstract
Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is the combined length of the Euclidean minimum spanning trees of red points and of blue points divided by the length of the Euclidean minimum spanning tree of the union of them. The maximum MST-ratio of a point set is the maximum MST-ratio over all non-trivial colorings of its points by red and blue. We prove that the problem of finding the maximum MST-ratio of a given point set is NP-hard when the dimension is a part of the input. Moreover, we present a ${O}(n^{2})$ running time $3$-approximation algorithm for it. As a part of the proof, we show that in any metric space, the maximum MST-ratio is smaller than $3$. Additionally, we study the average MST-ratio over all colorings of a set of $n$ points. We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ goes to infinity.
1 Introduction
Recently, motivated by applications in spatial biology, Cultrera et al. [6] studied the interactions between color classes in a colored point set from a topological point of view. They introduced the concept of MST-ratio as one of the measures for mingling points with different colors in a colored point set. Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees, or EMST for short, which is the main focus of this paper. For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, the Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is the minimum spanning tree of the complete geometric graph on $P$, where the edges have weights equal to the Euclidean distance between the two endpoints. For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the combined length of the Euclidean minimum spanning trees of $R$ and $B$, divided by the length of the Euclidean minimum spanning trees of $P$:
$\mu(P,R)=\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}.$
How much longer can the EMSTs of two finite sets be compared to the EMST of their union? Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitions of $P$ into two sets. Upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [7, 10], see Section 1.1 for further details. However, not a great deal of research has been conducted in the area of maximum EMST-ratio for higher dimensional Euclidean space.
Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set has yet been unanswered. The question can also be raised in the abstract setting. Namely, in the MAX-MST-ratio problem, given a weighted graph $G$ with positive weights, we aim to find the maximum over all bi-partitions of $V(G)$ of the combined length of the minimum spanning trees for induced subgraphs divided by the minimum spanning tree of $G$. As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a finite point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find MAX-MST-ratio of the complete geometric graph on $P$, with Euclidean distances as weights of edges. When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio (e.g. $2$-MAX-EMST-ratio, $3$-MAX-EMST-ratio). We tackle these computational problems, along with establishing bounds on the maximum and average MST-ratio, by presenting the following results:
  • The MAX-MST-ratio problem is NP-hard and even hard to approximate. (Theorems 1,2)
  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3)
  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 4)
  • There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 5)
  • For every set of $n$ points in the Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 7)
  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 6)
1.1 Related Works
The maximum EMST-ratio is closely related to the Steiner ratio of the Euclidean space. Recall that the Steiner tree of a point set is a tree that connects them via segments of minimum total length, allowing using some extra points. The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of the minimum spanning tree of the set. Figure 1 shows an example of three points forming an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$. Gilbert and Pollack [12] conjectured that this is the best example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$. The best-known lower bound for $\rho_{2}$ is $0.824$, due to Chung and Graham [5]. Important to us is also a universal lower bound on $\rho_{d}$ that works for every $d$. Gilbert and Pollack [12] presented a short proof for $\rho_{d}\geq 0.5$. Later, Graham and Hwang [13] showed that for any dimension $d$ we have $\rho_{d}\geq 0.577...$, which means the Steiner tree of a point set $P$ can never be shorter than 0.577×0.577..$\times$0.577 … × —EMST(P)—.
Figure 1: The point set $P$ shown by square nodes is just the three vertices of an equilateral triangle of side length 1. The black edges form EMST(P) with length $2$. Adding an extra vertex in the center, the green edges form the minimum Steiner tree of $P$ with length $\sqrt{3}$.
Using the above bounds for $\rho_{2}$, Cultrera et al. [7] showed that the supremum over all point sets in ${{\mathbb{R}}}^{2}$ of the maximum EMST-ratio is between $2.154$ and $2.427$. Note that it is not hard to see that the infimum of the maximum EMST-ratio is $1$. Their attention then shifted to lattice point sets, where they introduced the MST-ratio (despite the infinite number of points). They subsequently demonstrated that the infimum and supremum of the maximum MST-ratios across all $2$-dimensional lattices are $1.25$ and $2$, respectively.
Dumitrescu et al. [10] proved that for any set $P$ of at least $12$ points in ${{\mathbb{R}}}^{2}$, $\gamma(P)>1$. They also showed that for $n$ points sampled uniformly at random in $[0,1]^{2}$, the maximum EMST-ratio is at least $\sqrt{2}-{\varepsilon}$, for every ${\varepsilon}>0$, with a probability that tends to $1$ as $n$ goes to infinity. A tighter analysis shows that in the $d$-dimensional unit cube, the expected value of the average EMST-ratio over all colorings tends to $d]{2}$ as $n$ goes to infinity, see Theorem 6. The proof relies on the classic result by Beardwood, Halton, and Hammersley [3], saying that the length of a minimum spanning tree of a set $P$ of $n$ points uniformly distributed in $[0,1]^{d}$ satisfies
$|EMST(P)|/n^{1-1/d}$ $\rightarrow{}\beta(d)$ (1)
with probability one, where $\beta(d)>0$ is a constant depending only on the dimension. Best known bounds for $\beta(2)$ are approximately $0.6$ as a lower bound by Avram and Bertsimas [2], and $0.707$ as an upper bound by Gilbert [11].
1.2 Outline
Section 2 focuses on the abstract graphs and presents complexity results on the MAX-MST-ratio problem. Section 3 addresses the geometric setting and proves bound for the maximum MST-ratio as well as complexity results for MAX-EMST-ratio. Section 4 explores the average EMST-ratio over all colorings of a point set. Section 5 presents the experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets. Finally, Section 6 concludes the paper by presenting specific open questions aimed at deepening the understanding of the MST-ratio.
2 Maximum MST-ratio of Abstract Graphs
In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs. As the primary result of this section, we show that MAX-MST-ratio is NP-hard.
Theorem 1.
The problem MAX-MST-ratio is NP-hard.
Proof.
Let $\mathcal{I}$ be an instance of the well-known NP-complete $3$-SAT problem (see [8]), with a set of variables $X$ and set of clauses $\mathcal{C}$. Let $m=|\mathcal{C}|$ be the number of clauses of $\mathcal{I}$. Without loss of generality, we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$. Indeed, otherwise, one can deterministically decide the assignment of those variables and reach a smaller instance. Similarly, we can assume that in a fixed clause $C$, there is no variable $x$ that appears as $x$ and $\bar{x}$ as otherwise one can remove $C$ and again reach a smaller instance.
Using $\mathcal{I}$ we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
  • We introduce a new vertex $C$.
  • For every variable $x\in X$ that belongs to $C$ we add a new vertex $x_{C}$.
  • For every variable $x\in X$ such that $\bar{x}$ is in $C$ we add a new vertex $\bar{x}_{C}$.
  • Let $S_{C}$ be the vertex set that contains $C$ and all the vertices $x_{C},\bar{x}_{C}$ for every variable $x\in X$. Note that $|S_{C}|=4$. For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (purple edges in Figure 2).
For every variable $x\in X$, we define the variable gadget for $x$ in this way:
  • For every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$ then we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (orange edges in Figure 2).
We also add a vertex $F$ in our construction. For every clause $C$ we set the weight of the edge $FC$ to $1$. All the other edges in this graph have a weight $2$.
[…]
Figure 2: (a), (b), and (c) show the gadget for clauses $C_{1}=x\lor\bar{y}\lor z$, $C_{2}=\bar{x}\lor\bar{y}\lor z$, $C_{3}=\bar{x}\lor\bar{y}\lor\bar{z}$, respectively. (d) depicts the gadget of a variable $x\in X$. All the edges in in the figure have weight $1$.
In $\mathcal{I^{\prime}}$, given a bi-partition (i.e. a coloring) of vertices into sets $B$ and $R$, we are interested in $|MST(B)|+|MST(R)|$. Hence we refer to this as the value of this bi-partition (or coloring). We also denote the maximum possible value among all the possible colorings for $\mathcal{I^{\prime}}$ by $OPT(\mathcal{I^{\prime}})$. We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of the MST of $\mathcal{I^{\prime}}$.
Let $H$ be the induced subgraph of $\mathcal{I^{\prime}}$ on the edges of weight $1$. Observe that $H$ contains $m$ vertex-disjoint cliques of size four, which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$. We show that $\mathcal{I}$ is satisfiable iff $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Lemma 1.
If $\mathcal{I}$ is satisfiable, then $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Proof.
Let $A$ be an assignment that satisfies $\mathcal{I}$. We set the color of $F$ to blue and for every clause $C$ we set the color of the vertex $C$ to red. Now for every variable $x$:
  • if $A$ sets $x$ to true, then for every clause $C$ that contains $x$, we set the color of $x_{C}$ to blue and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to red.
  • otherwise, for every clause $C$ that contains $x$, we set the color of $x_{C}$ to red and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to blue.
Now we argue that the value of such coloring is $26m-$ 2. Firstly, observe that by construction, if the two endpoints of an edge of weight $1$ have the same color then they must belong to the same clause gadget. Therefore the subgraph induced by blue (and similarly red) vertices can only have edges of weight $2$ and the edges that belong to a clause gadget. Furthermore, since $A$ satisfies $\mathcal{I}$ then for every clause $C$ at least one of the vertices in $H_{C}$ is blue and by construction the vertex $C$ is colored in red. Altogether, if we consider the maximum forest on blue vertices and the maximum forest on red vertices of $H_{C}$, they have precisely two edges in total. Therefore, $MST(R)\cup MST(B)$ contains exactly $2m$ edges of weight $1$. Now since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ have $14m-$ 1 edges in total. Therefore $MST(R)\cup MST(B)$ has exactly $(4m-1)-2m=2m-1$ edges of weight $2$ and hence we have $|MST(R)|+|MST(B)|=6m-2$ and hence the claim. ∎
Lemma 2.
If $OPT(\mathcal{I^{\prime}})\geq 6m-2$, then $\mathcal{I}$ is satisfiable.
Proof.
Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of blue and red vertices then $MST(R)+MST(B)\geq 6m-2$. First, assume that there is an edge $uv$ of weight $1$ such that $u$ and $v$ are colored with the same color in $Q$ but do not belong to the same clause gadget. Then $MST(R)\cup MST(B)$ contains at least $12m+$ 1 edges of weight $1$, as $H$ contains $m$ disjoint cliques of size $4$ and also the edge $uv$. Since $MST(R)\cup MST(B)$ has $14m-$ 1 edges then $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction. This means that any edge of weight $1$ whose endpoints have the same color belongs to a clause gadget.
Now let us construct an assignment $A(Q)$, that satisfies $\mathcal{I}$. Without loss of generality, assume that $Q$ colors the vertex $F$ in blue. Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red. Also, by the above argument for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ must be different in $Q$. Now consider a variable $x$. $A(Q)$ sets $x\leftarrow TRUE$ if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by red or it colors $\bar{x}_{C}$ by blue. We claim that every clause in $\mathcal{I^{\prime}}$ is satisfied. Assume this is not the case and a clause $C$ is not satisfied. This implies that all the vertices in $H_{C}$ are colored in red. Now if we take any other clause $C^{\prime}$, in $MST(R)\cup MST(B)$ there are at least two edges from $H_{C^{\prime}}$. This means that $MST(R)\cup MST(B)$ contains at least $13+2(m-1)=2m+$ 1 edges of weight $1$. Thus $|MST(R)|+|MST(B)|\leq(2m+1)+(4m-1-(2m+1))\times 2<6m-2$, a contradiction. ∎
This completes the proof and hence we can conclude that computing MAX-MST-ratio is NP-hard. ∎
In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard. Moreover, in the general setting, the MAX-MST-ratio problem is extremely difficult to approximate. Theorem 2 presents a strong inapproximability result for computing the MAX-MST-ratio.
Theorem 2.
For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio, unless P=NP.
Proof.
We first prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem.
Let $G=(V,E)$ be a graph on $n$ vertices on which we wish to find the maximum clique. From this we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and for each edge $e\in E(G^{\prime})$ its weight is $n$ if $e\in E$, and it is one otherwise.
First, we claim that given a clique $C$ of size $\ell$ in $G$, we can find a coloring of value at least $(\ell-2)n+(n-\ell)$ for $G^{\prime}$ in polynomial time. This is trivial if $\ell\leq 2$, and assuming $\ell\geq 3$ it is sufficient to assign the color red to all but one of the vertices in $C$, and blue to the remaining vertices. In this case, the sum of the weights in the MST of the red vertices is $(\ell-2)n$ and for the blue vertices the MST has exactly $(n-\ell)$ many edges and hence the claim follows. Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lceil k/n\rceil}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ be the red MST and blue MST of $G$ in the coloring $Q$. Note that given the weights in $G$, and since $R_{Q}$ and $B_{Q}$ have $n-2$ edges in total, then in total $R_{Q}$ and $B_{Q}$ have exactly $\lceil k/n\rceil$ edges of weight $n$. Without loss of generality, we can assume that $R_{Q}$ has at least $\frac{\lceil k/n\rceil}{2}$ edges of weight. Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight one in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, $r\geq\frac{\lceil k/n\rceil}{2}+1$. For every $1\leq i\leq k$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is the MST on the set of red vertices then for every $i<j$, the weight of the edge $u_{i}u_{j}$ is $n$. Hence $\{u_{1},u_{2},\ldots,u_{k}\}$ is a clique of size $\frac{\ell}{2}+1$ in $G$.
Altogether, this shows that one can use an $\alpha$-approximation algorithm for the MAX-MST-ratio problem to find a $2\alpha$-approximation for the MAX-Clique problem. Moreover, Zuckerman [17] proved that there exists no $O(n^{1-\varepsilon})$-approximation for MAX-Clique problem unless $P=NP$. Therefore, there exists no $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio. ∎
3 Maximum MST-ratio of Geometric Point Sets
A weighted complete graph $G$ on $n$ vertices is realizable in $\mathbb{R}^{d}$ if there exist $n$ points in $\mathbb{R}^{d}$ such that the distance graph of these points is isomorphic to $G$.
We rely on a result by Dekster and Wilker [9], which shows that roughly speaking, any weighted complete graph with positive weights that do not differ too much, is realizable in Euclidean space.
Lemma 3.
[9, Theorem 2] For every positive integer $n>0$ there exists a real number $10<\lambda_{n}<$ 1 such that every complete graph $G$ on $n$ vertices whose all edges have weights in $[\lambda_{n}\ell,\ell]$, for a positive real number $\ell$ is realizable in ${{\mathbb{R}}}^{n-1}$.
We use the above lemma to increase all the weights of an instance in Theorem 1 by a fixed number in order to make it realizable in Euclidean space.
Lemma 4.
For every weighted complete graph $G$ on $n$ vertices, there exists a number $N>0$ such that if we increase the weight of every edge in $G$ by $N$, then the resulting graph is realizable in ${{\mathbb{R}}}^{n-1}$.
Proof.
Let $k=\binom{n}{2}$ and let $w_{1}\leq w_{2}\leq\ldots\leq w_{k}$ be the weights of the edges of $G$. Pick $N$ large enough such that $N>|w_{1}|$ and $N>\frac{\lambda_{n}w_{k}-w_{1}}{1-\lambda_{n}}$, where $\lambda_{n}$ is the real number taken from Lemma 3. Increasing the weights of the edges in $G$ by $N$ results in a graph $G^{\prime}$ with positive weights $N+w_{1}\leq N+w_{2}\leq\ldots\leq N+w_{k}$. $N$ being large enough guarantees $N+w_{1}\geq\lambda_{n}(N+w_{k})$. So $G^{\prime}$ satisfies the conditions of Lemma 3, and therefore is realizable in ${{\mathbb{R}}}^{n-1}$. ∎
Theorem 3.
The problem MAX-EMST-ratio is NP-hard.
Proof.
By Lemma 4, the constructed instances in Theorem 1 can be turned into a geometric graph by increasing the weights of all edges by a fixed amount. Note that this action does not change the optimal coloring of the graph in terms of the maximum MST-ratio. This is because the total number of edges in the minimum spanning trees of the two colors always falls two short of the total number of vertices. Thus, from any instance of $3$-SAT problem, we can construct an instance of the MAX-EMST-ratio problem. ∎
As established in Theorem 3, determining the maximum EMST-ratio is NP-hard. We conjecture that the problem remains NP-hard even when restricted to a specific dimension. In particular, we are interested in points in ${{\mathbb{R}}}^{2}$ and ${{\mathbb{R}}}^{3}$, as they are relevant for practical applications.
Conjecture 1.
The problems $2$-MAX-EMST-ratio and $3$-MAX-EMST are NP-hard.
Theorem 2 shows that it is even hard to approximate the maximum MST-ratio for general graphs. However, in Euclidean space, it is not difficult to approximate the maximum EMST-ratio as it cannot be too large. We show a general upper bound for the maximum MST-ratio in any metric space that leads to a simple $3$-approximation algorithm for the MAX-EMST-ratio problem. In the proof, we use the following well-known fact about the path double cover of trees.
Lemma 5 ([15]).
Given a tree $T$ with $\ell$ leaves, there exist $\ell$ (not necessarily pairwise distinct) paths in $T$ such that each edge of $T$ belongs to precisely two of these paths.
Theorem 4.
In any metric space, the maximum MST-ratio for every set of $n$ points ($n\geq 5$) is not more than $3-\frac{4}{n-1}$.
Proof.
Consider an arbitrary point set $S$ in a metric space, and let $T$ be a minimum spanning tree of $S$. To prove the statement, for every coloring $Q$ of $S$ we build spanning trees $R_{Q}$ on the set of red vertices, and $B_{Q}$ on the set of blue vertices of $Q$, such that the total weight of $R_{Q}\cup B_{Q}$ is at most $(3-\frac{4}{n-1})|T|$. Let $P^{*}$ be a path with maximum total weight among all paths in $T$ whose endpoints are not leaves of $T$. Assume, without loss of generality, that one endpoint of $P^{*}$, denoted by $r$, is red, and we hang $T$ from $r$. We traverse the tree from the root $r$ using Depth-First Search (DFS), prioritizing the vertices of $P^{*}$ when exploring branches. The starting time of a node refers to the moment it is first visited during the traversal.
For an arbitrary coloring $Q$ of $S$, we build $R_{Q}$ and $B_{Q}$ in the following way (see Figure 3):
  • We add every edge $e$ of $T$ that has both endpoints in red (in blue) to $R_{Q}$ (to $B_{Q}$).
  • Consider any red (blue) vertex $v$ in $T$ and any blue (red) child $w$ of $v$ such that there exists at least one red (blue) vertex in the descendants of $w$ (e.g. $v=r$ in Figure 3). Let $u_{1},u_{2},\ldots,u_{x}$ be the red descendants of $w$, sorted by starting time (i.e. $u_{1}$ has the earliest starting time, and $u_{x}$ the latest), for which all the internal vertices of the $wu_{i}$-path in $T$ are blue (red). We add the edges $u_{1}u_{2},u_{2}u_{3},\ldots,u_{x-1}u_{x},u_{x}v$ to $R_{Q}$ (to $B_{Q}$).
  • Let $v_{1}$, $v_{2}$, $\ldots$, $v_{y}$ be the set of blue vertices that have no blue ancestors, sorted by starting time. We add the edges $v_{1}v_{2},v_{2}v_{3},\ldots,v_{y-1}v_{y}$ to $B_{Q}$.
$r$$u_{3}$$u_{2}$$u_{1}$$w$
Figure 3: The construction of $R_{Q}$ and $B_{Q}$ (dashed edges) according to the coloring $Q$ of vertices in the proof of Theorem 4. The solid black and green edges form the tree $T$ rooted at $r$. The children of any vertex are from left to right according to the order of their starting time in DFS. The green edges are the edges of $P^{*}$.
Let $\mathcal{P}$ be the set of all $uv$-paths $P_{e}$ in $T$, where $e=uv$ is an edge in $R_{Q}\cup B_{Q}$. Observe that every edge of $T$ belongs to at most $3$ paths in $\mathcal{P}$. Namely, by construction, if $x$ is the parent of $y$ and $x$ is blue (similarly red) then the edge $xy$ appears in at most one path $P_{e_{b}}$ where $e_{b}\in B_{Q}$ and at most two paths $P_{e_{r}}$ where $e_{r}\in R_{Q}$. Furthermore, there are certain edges in $T$ that are covered by at most two paths in $\mathcal{P}$, namely the leaf edges (i.e. edges incident to a leaf) and the edges in $P^{*}$ (i.e. the green path in Figure 3).
Since the instance is metric, the weight of every edge $uv$ in $|R_{Q}|\cup|B_{Q}|$ is not more than the total weight of the corresponding $uv$-path in $\mathcal{P}$. So we have $|R_{Q}|+|B_{Q}|\leq 3|T|-W_{L}-W_{P^{*}}$, where $L$ is the set of leaf edges in $T$ and $W_{L}$ and $W_{P^{*}}$ are the total weight of the edges of $L$ and $P^{*}$, respectively. It suffices to show that $W_{P^{*}}+W_{L}\geq(\frac{4}{n-1})|T|$.
If the diameter of $T$ is less than $4$, then $P^{*}\cup L=T$, and hence the claim follows as $W_{P^{*}}+W_{L}=|T|\geq(\frac{4}{n-1})|T|$. Assume that the diameter of $T$ is at least $4$. Let $T^{\prime}$ be the tree obtained from $T$ by removing all the leaves of $T$ and let $\ell$ and $\ell^{\prime}$ be the number of leaves of $T$ and $T^{\prime}$, respectively. Note that a vertex in $T$ that is a leaf in $T^{\prime}$ must be adjacent to at least one leaf in $T$ and hence $\ell^{\prime}\leq\ell$. Furthermore as the diameter of $T$ is at least $4$, then $T^{\prime}$ has at least one non-leaf vertex which implies that $\ell+\ell^{\prime}\leq n-1$ and thus $\ell^{\prime}\leq\frac{n-1}{2}$.
Using Lemma 5, there are $\ell^{\prime}$ paths in $T^{\prime}$ such that all edges of $T^{\prime}$ belong to exactly two such paths. Hence $W_{P^{*}}$ is lower bounded by $\frac{2|T^{\prime}|}{\ell^{\prime}}=\frac{2(|T|-W_{L})}{\ell^{\prime}}\geq\frac{4(|T|-W_{L})}{n-1}$.
Therefore $W_{P^{*}}+W_{L}\geq\frac{4(|T|-W_{L})}{n-1}+W_{L}\geq\frac{4}{n-1}|T|.$
Note that the above bound is the best we can achieve in metric spaces. In fact, for every odd $n$, there exists a metric space with $n$ points (elements) for which the maximum MST-ratio is $3-\frac{4}{n-1}$. To see this, let $n=2k+1$, and define the points $v_{1},\ldots,v_{n}$. For every $i<n$, we set the weight of $v_{n}v_{i}$ to be one. For every $1\leq i\leq k$, we set the weight of $v_{2i}v_{2i-1}$ to be zero. We set the weight of any other edge to two. We first argue that the weights are metric. Note that the set of possible weights are $\{0,1,2\}$, the edges of weight one are the edges incident on $v_{n}$, and the edges of weight zero form a matching. Therefore, any triangle with edge weights $a\leq b\leq c$ has either zero or two edges of weight one and has at most one edge of weight zero. Thus $(a,b,c)$ is either $(0,1,1)$, or $(0,2,2)$, or $(1,1,2)$, or $(2,2,2)$, in any case we have $c\leq a+b$. Observe that $|MST(S)|=0\times k+1\times k=k$. If we color $v_{1},v_{3},\ldots,v_{n}$ by red all the other vertices by blue, then $|MST(R)|=k$ and $|MST(B)|=2(k-1)=2k-2$. Therefore $|MST(R)|+|MST(B)|=3k-2$ and hence the MST-ratio is $\frac{3k-2}{k}=3-\frac{2}{k}=3-\frac{4}{n-1}$.
Theorem 5.
There exists a ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem.
Proof.
By Theorem 4, the maximum EMST-ratio for an $n$-element point set is bounded from above by $3-\frac{4}{n-1}$. On the other hand, one can achieve EMST-ratio of at least $\frac{n-2}{n-1}$ by first computing the minimum spanning tree of the point set, and then coloring the vertices of the two components of it after removing the shortest edge by red and blue. The running time for computing this coloring is ${O}(n^{2})$, see [8, 16], and clearly, $\frac{n-2}{n-1}\geq\frac{1}{3}.(3-\frac{4}{n-1})$. ∎
We remark that one could achieve a weaker approximation factor $3.47$ in a simpler way, by using a trivial upper bound $\frac{2}{\rho_{d}}\leq\frac{2}{0.577...}<3.47$ on the MAX-EMST-ratio.
Indeed, if we restrict to ${{\mathbb{R}}}^{2}$ (i.e. $2$-MAX-EMST-ratio), the approximation factor in Theorem 5 can be improved to $2.427+\varepsilon$, because the upper bound for the maximum EMST-ratio in the plane is $\approx 2.427$ (see [7]). As can be observed, the coloring used to present the $3$-approximation is fairly straightforward. However, there seems to be potential for improvement by discovering more refined colorings in terms of mixedness.
We introduce the Bipartite Coloring as a good candidate to approximate the maximum EMST-ratio of a point set. By bipartite coloring, we mean the partition of a point set into two subsets labeled by colors in which every edge of the minimum spanning tree has different colors on the endpoints. Intuitively, it gives a well-distributed coloring of the point set that can be computed in ${O}(n^{2})$ time. Our experiments suggest that the bipartite coloring gives a very good approximation of the maximum EMST-ratio for random point sets in the plane, see Section 5. Although the bipartite coloring seems to be effective, it does not always produce the maximum EMST-ratio. Here is an extreme example:
Consider $12n+$ 1 points on a zigzag path in the triangular grid and call them a Triangular Chain. Figure 4 illustrates a triangular chain for $n=5$. The maximum EMST-ratio of a triangular chain with $12n+$ 1 points is $\frac{3n-2}{2n}$ while the ratio for the bipartite coloring is $\frac{2n-1}{2n}$. (By stretching the chain slightly, one can ensure a unique MST and, consequently, a unique bipartite coloring.) As $n$ increases, the proportion between the maximum EMST-ratio and the bipartite EMST-ratio approaches $1.5$.
Figure 4: An example of a point set where the maximum EMST-ratio has a significant gap from the EMST-ratio of the bipartite coloring. The left panel shows a triangular chain for $n=5$, its EMST in green, and the red and blue EMSTs in the bipartite coloring. The right panel presents the coloring that achieves the maximum EMST-ratio.
We conjecture that this is the worst approximation factor of the maximum EMST-ratio that can be obtained from bipartite coloring for a point set in the plane. Recall that if this conjecture is correct, it would lead to a $O(n^{2})$ time algorithm that achieves a $(1.5)$-approximation for the $2$-MAX-EMST-ratio problem.
Conjecture 2.
For every point set in ${{\mathbb{R}}}^{2}$, the EMST-ratio for the bipartite coloring is a $(1.5)$-approximation for the maximum EMST-ratio.
Apart from the bipartite coloring, we are interested in any improved approximation algorithm for the MAX-EXST-ratio that can be obtained by a simple coloring of a point set.
Question 1.
Is there any polynomial time algorithm for $2$-approximation of the MAX-EMST-ratio problem?
4 Average MST-ratio.
The Average EMST-ratio of a point set $P$ is the average over all $2$-colorings of $P$ of the EMST-ratio. It is automatically a lower bound for the maximum MST-ratio. Besides, it could be considered as a measure to see how close a point set is to a typical random point set, as we know the behavior of random point sets. In fact, due to Ondřej Draganov and Morteza Saghafian, the expected value for the average EMST-ratio of a random point set in a unit square is $\sqrt{2}$ in the limit. Here, we present the proof for the generalized version of this result, meaning for points within a $d$-dimensional unit cube.
Theorem 6.
For a set of $n$ random points uniformly distributed in $[0,1]^{d}$, the expected value of the average EMST-ratio tends to $d]{2}$ as $n$ goes to infinity.
Proof.
Let $n$ be large enough and $P$ be a set of $n$ random points uniformly distributed in $[0,1]^{d}$. Consider all $22^{n}-$ 2 proper bipartitions $P=B\cup R$ into blue and red points. Clearly, for $11\leq k\leq n-$ 1, in $\binom{n}{k}$ of such bipartitions $|B|=k,|R|=n-k$. Ignoring the relatively small $k$ or $n-k$, by (1), we have
$|EMST(B)|\sim\beta(d).k^{1-1/d},\hskip 14.22636pt|EMST(R)|\sim\beta(d).(n-k)^{1-1/d},\hskip 14.22636pt|EMST(P)|\sim\beta(d).n^{1-1/d}$
with probability one, where $\beta(d)$ is a constant. Thus, the average EMST-ratio over all proper bipartitions of $P$ is
$\sum_{P=B\cup R}\frac{|EMST(B)|+|EMST(R)|}{|EMST(P)|}$ $\sim\sum_{k=1}^{n-1}\frac{[\beta(d).k^{1-1/d}+\beta(d).(n-k)^{1-1/d}].\binom{n}{k}}{\beta(d).n^{1-1/d}.(2^{n}-2)}$
$=\sum_{k=1}^{n-1}\frac{[k^{1-1/d}+(n-k)^{1-1/d}].\binom{n}{k}}{n^{1-1/d}.(2^{n}-2)}$
with probability one as $n$ goes to infinity. The right-hand side is a special case of convergence of Bernstein polynomials (see [4, 14]), and therefore converges to $d]{2}$ as $n$ goes to infinity. ∎
Certainly, Theorem 4 implies that $3$ is an upper bound for the average EMST-ratio of point sets in the Euclidean space. How much can this upper bound be improved? There exist point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio $\approx 2.154$.
[…]
Figure 5: An example of a point set with a large average EMST-ratio consists of the origin and several points clustered near the three vertices of an equilateral triangle centered at the origin. As the number of points increases, the average EMST-ratio approaches approximately $2.154$.
For instance, consider the unit circle through the origin in the plane, an equilateral triangle $abc$ on it, and then a point set $P=A^{\prime}\cup B^{\prime}\cup C^{\prime}\cup\{0\}$, where $A^{\prime}$ (resp. $B^{\prime}$, $C^{\prime}$) consists of many points that lie within distance $\varepsilon$ of $a$ (resp. $b$,$c$), for $\varepsilon$ being small enough, see Figure 5. In this case, the average EMST-ratio approaches $\frac{3+2\sqrt{3}}{3}\approx 2.154$ as the number of points increases, because with high probability, in a random two coloring, each of the sets $A^{\prime}$, $B^{\prime}$ and $C^{\prime}$ will contain both colors.
As for the lower bound, we show that the average EMST-ratio for $n$ points is always not less than $\frac{n-2}{n-1}$. This is a stronger result than the first part of Theorem 2 in [10].
Theorem 7.
For every set of $n$ points in the Euclidean space, the average EMST-ratio is at least $\frac{n-2}{n-1}$.
Proof.
Consider a set of $n$ points $P$ and let $e_{1},...,e_{n-1}$ be the edges in an Euclidean minimum spanning tree $T$ of $P$. Let us assume that the edge $e_{i}$ has weight $w_{i}$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$. Considering a coloring of $P$ in which $R$ is the set of red points and $B=P\setminus R$ is the set of blue points, we call an edge of $T$ colorful if the two endpoints of it have different colors. Let $i$ be the smallest index such that the edge $e_{i}$ is colorful. $EMST(R)\cup EMST(B)\cup\{e_{i}\}$ is a spanning tree of $P$ and therefore, its weight is at least $|EMST(P)|$. This implies $|EMST(R)|+|EMST(B)|\geq|EMST(P)|-w_{i}$.
On the other hand, considering a random coloring of $P$ with two colors, the probability for the edge $e_{i}$ being colorful is $\frac{1}{2}$, and since the $e_{i}$’s do not form a cycle, their corresponding events are independent. Thus, for a fixed index $j$, the probability that in a random coloring, the shortest colorful edge of $T$ is $e_{j}$ is $(\frac{1}{2})^{j}$. Therefore, there are precisely $2^{n-j}$ colorings of $P$ in which $j$ is the smallest index such that the edge $e_{j}$ is colorful.
By the above arguments and since we have $22^{n}-$ 2 possible (proper) colorings, the average EMST-ratio is at least:
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}.$
As $\sum_{i=1}^{n-1}2^{n-i}=2^{n}-2$ and $w_{1}\leq w_{2}\leq\ldots\leq w_{n-1}$, then:
$|MST(P)|-\sum_{i=1}^{n-1}\frac{2^{n-i}w_{i}}{2^{n}-2}\geq|MST(P)|-\frac{|MST(P)|}{n-1},$
and hence the claim. ∎
Note that Theorem 7 holds also for general graphs.
The average EMST-ratio appears to be greater than 1 in many cases, though this is not always true. The examples we found in ${{\mathbb{R}}}^{2}$ consist of point sets where $n-5\leq k\leq n-1$ points are clustered closely together, forming a core, and $1\leq n-k\leq 5$ points are sufficiently far from the core such that their pairwise distances exceed their distances to the core, see Figure 6. We believe those are the only point sets with an average EMST-ratio less than $1$.
Figure 6: A point set with average EMST-ratio smaller than $1$.
Question 2.
Can the point sets in ${{\mathbb{R}}}^{2}$ with average EMST-ratio less than $1$ be characterized?
5 Computational Experiments
This section presents the results of computational experiments comparing the average EMST-ratio, maximum EMST-ratio, and the EMST-ratio for bipartite coloring, all conducted on random point sets. The setup is as follows. In a single experiment, we consider $n$ points selected uniformly at random in the unit square, we compute the EMST-ratios for all $22^{n}-$ 2 proper colorings of this point set, and we take the maximum and the average over all colorings. Additionally, we compute the EMST-ratio for the bipartite coloring. For $n$ between $5$ and $20$, we compute the mean values of the functions (i.e. the average EMST-ratio, maximum EMST-ratio, and bipartite EMST-ratio) across 100 experiments. The resulting averages are displayed in Figure 7. According to Theorem 6, the average EMST-ratio (the red curve) tends to $\sqrt{2}$. Interestingly, the diagram in Figure 7 suggests that a similar behavior might hold for the maximum EMST-ratio (the green curve) as well, though approaching a different limit. In fact, the program did not find any instance of a random point set with a maximum MST-ratio greater than $2$ in our experiments.
Figure 7: The mean values of the maximum EMST-ratios (green), the bipartite EMST-ratios (blue), and the average EMST-ratios (red) of $100$ random sets of $n$ points, where $5\leq n\leq 20$.
Question 3.
What is the limit for the expected maximum EMST-ratio of $n$ points sampled uniformly at random in $[0,1]^{2}$, as $n$ goes to infinity?
Conjecture 3.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $2$ with probability tending to $1$ as $n\rightarrow{\infty}$.
As for the bipartite coloring, Figure 7 clearly shows how close its EMST-ratio is to the maximum EMST-ratio in expectation. Figure 8 particularly compares the EMST-ratio for the bipartite coloring and the maximum EMST-ratio across $100000$ experiments involving $5$ to $20$ points. Moreover, the maximum EMST-ratio is never more than $1.3$ times and is rarely more than $1.1$ times the bipartite EMST-ratio across all these experiments.
Figure 8: Each point in the diagram represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the maximum EMST-ratio, respectively.
Conjecture 4.
For $n$ random points uniformly distributed in $[0,1]^{2}$, the maximum EMST-ratio is less than $1.1$ times the bipartite EMST-ratio with probability tending to $1$ as $n$ goes to infinity.
Figure 9: In this diagram each point represents a random point cloud, with the $x-$ and $y-$coordinates corresponding to the bipartite EMST-ratio and the average EMST-ratio, respectively.
Figure 9, on the other hand, presents a comparison between the EMST-ratio for the bipartite coloring and the average EMST-ratio obtained from $100000$ experiments with point sets ranging from $5$ to $20$ points. Interestingly, in all computed experiments without exception, the bipartite EMST-ratio exceeds the average EMST-ratio. This suggests that the bipartite coloring of a point set provides a better EMST-ratio than what is expected from a random coloring. Although this is not always the case for finite point sets, we believe there are very few counterexamples.
Question 4.
Can the point sets in ${{\mathbb{R}}}^{2}$ for which the bipartite EMST-ratio is smaller than the average EMST-ratio be characterized?
6 Discussion
In addition to the conjectures and questions posed in this paper, there are several other directions worth exploring further.
Coloring with Many Crossings.
For a point set $P$ in the plane and a partition $P=R\cup B$ of $P$ into red and blue points, the Chromatic Crossing Number is defined as the number of crossings between pairs of edges, one in $EMST(R)$ and one in $EMST(B)$. Intuitively, this number can be viewed as another measure of how mixed the coloring is; the higher the chromatic crossing number, the more mingled the coloring. In fact, for any point set, it is easy to find colorings with no chromatic crossings. More interesting, however, is the Maximum Chromatic Crossing Number, which represents the highest chromatic crossing number achievable across all possible colorings of the point set.
Question 5.
How large and how small can the maximum chromatic crossing number of a set of $n$ points in the plane be?
Question 6.
Is there any relation between the chromatic crossing number and the EMST-ratio? In particular, does the coloring that maximizes the EMST-ratio tend to have a high chromatic crossing number?
Conjecture 5.
Given a point set $P$ in the plane, determining the maximum chromatic crossing number of $P$ is NP-hard.
Maximum MST Length over Subsets.
Related question to finding the maximum MST-ratio is the following. In the MAX-EMST-subset problem, given a point set $P$, we aim to find the maximum EMST length over all subsets $Q\subseteq P$. The question can be raised also in the abstract setting. Namely, in the MAX-MST-subset problem, given a weighted graph $G$, we would like to find the maximum MST length over all induced subgraphs $H\subseteq G$.
Question 7.
What can we say about the complexity of MAX-MST-subset and MAX-EMST-subset problems? Are they NP-hard?
More Colors.
Similar questions can be asked when we color a point set with $c>2$ colors. To express the MST-ratio, we consider the combined length of the minimum spanning trees of points in each color divided by the length of the minimum spanning tree of the (uncolored) point set. Theorem 4 can be generalized to obtain $12c-$ 1 as an upper bound for the MST-ratio in any metric space.
Acknowledgements. The authors thank Ondřej Draganov for his valuable insights on the topic of average EMST-ratio.